### v1

[…]
H1.
Most of its components are 3D-printed using polylactic acid (PLA), leading to low rigidity and susceptibility to structural deformation.
H2.
Unlike typical robots with constrained ranges, its joints can rotate beyond 360∘, potentially reaching positions the robot cannot achieve.
H3.
Although portable, inevitable body movements during operations can cause its base to shift, leading to inaccurate action recording.
H4.
Its gripper lacks smooth control, leading to potential jamming under clamping forces.
H5.
Its shaft connects directly to the encoder, with wires routed along its side. Joint movement can cause constant friction and stretching, risking wire breakage or short circuits over long-time use.
During demonstration collection using teleoperation, most of the above issues (except H4) might not significantly impact data acquisition, as the human operator can adjust their actions based on the movements of the robotic arms. However, during the in-the-wild demonstration collection, these drawbacks can substantially affect motion capture accuracy and may result in invalid or unusable demonstration data. The hardware design of AirExo-2 is primarily driven by the limitations mentioned above. To ensure seamless integration with the learning process from in-the-wild demonstrations, we design the exoskeleton to match the dimensions of the robotic arm in a 1:1 ratio. This design choice helps avoid unnecessary obstacles in direct learning from the collected demonstrations. The key hardware designs are outlined as follows. Enhanced Overall Structural Rigidity (H1). In AirExo, the links connecting two consecutive joints are the most prone to deformation. In AirExo-2, we replace the 3D-printed parts with 20x20 European standard aluminum profiles, providing significantly higher strength at a very low cost. For the joints, the outer shell is 3D-printed using PLA-CF, a carbon fiber-reinforced PLA material with higher hardness. Inside the joint, larger bearings are used to further enhance structural rigidity. Together with the improved links, the hardware upgrade significantly increases the overall structural rigidity of AirExo-2, making the exoskeleton more durable, and thereby improving data collection accuracy. Hollow Rotating Disc and Side-Mounted Encoder (H5). As shown in Fig. H1., this design features an encoder mounted on the side of the joint, which uses a gear mechanism to translate the rotational angle of the rotating disc into encoder readings. The hollow disc design allows wires to pass through, preventing them from stretching during rotation and thereby extending their lifespan. Additionally, the side-mounted encoder simplifies maintenance, enabling easy debugging and replacement without the need to disassemble the joint. Joint with Angle Limit and Adjustable Friction (H2). This structure consists of a rotating disc with a grooved track and a friction pad that can be embedded into the track. Together, they allow for adjustment of the limiting angle, ensuring that the motion range of the AirExo-2 joint exactly aligns with the corresponding joint range of the robot. As illustrated in Fig. H1., the rotating disc, pre-joint, and post-joint are connected through bearings, allowing the rotational motion of the joint to be directly transmitted to the track. For demonstration collection, excessive or insufficient friction in the rotation of the joints is undesirable. Hence, the friction force of the joint in AirExo-2 can be adjusted by turning the screw on the outer shell, which compresses the friction pad. This design ensures optimal friction for comfortable and accurate data collection.
[width=]images/hardware.pdf
Smooth Gripper Control (H4). Following [umi, forcemimic], the gripper of AirExo-2 incorporates a linear guide, with the fingers mounted on a sliding block. This design allows for smoother opening and closing of the gripper, ensuring it operates seamlessly even under significant clamping forces without any stalling. Mobile Data Collection Platform (H3). Portability is crucial for in-the-wild data collection. However, to address the issue of base movement caused by the body motion of the operator, we mount AirExo-2 on a mobile aluminum profile stand, as shown in Fig. H1.. This setup ensures stability of the base during demonstration collection while maintaining the flexibility needed for mobility, enabling large-scale demonstration collection in real-world environments. An Intel RealSense D415 camera is set up on the top of the mobile platform to capture global observations. We also designed two optional camera mounts (though not used in this paper) for the future integration of in-hand cameras on the top of both grippers.
[…]
The AirExo-2 system requires two types of calibration simultaneously: (1) aligning the zero positions of each joint with the corresponding robot joint, and (2) determining the transformation between the global camera and the AirExo-2 base. To address these challenges, we propose a two-stage calibration process. Initial Calibration. For initial calibration, the former calibration can be achieved by manually adjusting the joints to approximate the zero position using specialized 3D-printed tools and reading the encoder values, obtaining . The latter calibration can be done by attaching an ArUco calibration marker board with a known position on the base and performing optical calibration using the OpenCV library [opencv], obtaining . Thus, the transformation can calculated as
However, this approach introduces errors due to human observation, calibration board misalignment, and optical inaccuracies. In a chained system like AirExo-2, these errors may propagate and amplify across joints, leading to significant cumulative inaccuracies in the end-effector pose. Therefore, fine-grained calibration is essential to ensure precise and consistent alignment between AirExo-2 and the camera frames during demonstration collection. Calibration via Differentiable Rendering. In the second stage, inspired by prior works [easyhec, easyhec++, sam-rl], we use differentiable rendering [differentiable_rendering] to refine the initial calibration. Training samples are obtained from a single human play trajectory with AirExo-2. Using the joint states and calibration parameters, we render the system mask and depth via a differentiable rendering engine [redner]. Calibration parameters are optimized by minimizing discrepancies between the rendered and annotated system masks, as well as between the rendered and observed depths. Pseudo-ground-truth masks are manually annotated with SAM-2 [sam2]. This iterative refinement compensates for errors that accumulate across joints, ultimately improving the overall accuracy. Specifically, we define , the calibration parameters to be optimized through differentiable rendering, as
where parameters, except for the base-to-camera rotation, are represented as deltas relative to the initial calibration results. The base-to-camera rotation is expressed in a 6D format [rot6d]. Thus, the final calibration results can be calculated as
{gather}
[…]
and the initial parameter values are set as
For the optimization process, we first record a single in-the-wild play trajectory, in which the human operator uses the AirExo-2 to adopt various poses. During trajectory recording, ensure that all parts of the AirExo-2 remain above the human hands and arms from the camera’s perspective. After data collection, we sample approximately 40 image-joint pairs from the trajectory, denoted as , where represents the total number of training samples for calibration, and and are the RGB and depth images of the -th sample, respectively. Subsequently, we utilize SAM-2 [sam2] to annotate the AirExo-2 mask and the depth mask for each sample , as shown in Fig. H1.. The first mask provides supervision for the rendered AirExo-2 mask, while the second mask is used to select the valid depth information that serves as the supervision signal.
[width=]images/calib_detail.pdf
The differentiable rendering engine Redner [redner] is employed to render the AirExo-2 mask and AirExo-2 depth using the calibration results and joint information:
where the rendering engine computes the gradients of the calibration parameters during the rendering process, and represent the initial calibration results. The rendered AirExo-2 mask is supervised by the human-annotated pseudo-AirExo-2 mask , and the rendered AirExo-2 depth is supervised by the camera depth within the region of the human-annotated depth mask . The depth mask ensures that only accurate depth information contributes to the loss. Thus, the objective can be written as:
where represents the weighting coefficient, and denotes the mask-apply operation. In practice, we set , use samples for optimization, and employ the Adam optimizer [adam] with a learning rate of for 1000 iterations to fine-tune the calibration results.
[…]
Apart from the two-stage calibration process described in Appendix H1., we implement several calibration alternatives, including (1) initial calibration, which uses the calibration results from the first stage without further fine-tuning; (2) human annotation, where a human operator utilizes a real-time GUI program to manually adjust the calibration parameters to align the rendered AirExo-2 contour with the visually observed contour from camera frames; and (3) two-stage calibration (mask only), which fine-tunes the calibration results by using only mask differences as supervision.
[…]
| Human Annotation | ||
| Two-Stage Calibration (mask only) | ||
| Two-Stage Calibration (mask + depth) | ||
[…]
Our two-stage calibration process achieves more accurate results compared to other alternatives. As reported in Tab. H1., our two-stage calibration process achieves the lowest error rates, with a 0.78% mask difference and 14.0 mm depth difference, yielding more precise calibration. It’s worth noticing that the depth difference here does not fully represent action accuracy because commercial depth sensors can produce noisy depth maps. Please refer to §LABEL:sec:accuracy-anaylsis for more details about action accuracy. Interestingly, human annotation performs even worse than initial calibration, mainly because annotators can only rely on 2D visual information to adjust calibration parameters. This limitation makes it difficult to accurately estimate depth information, leading to larger errors. Conversely, our two-stage calibration process explicitly models 3D information via differentiable rendering, offering a more reliable and precise solution for calibrating the AirExo-2 system.
[…]
Semi-Automatic SAM-2 Annotations. In the image adaptor, we initially annotate the hand mask (and, if visible, the head mask) manually using SAM-2 [sam2]. However, after a few annotations, we can fine-tune SAM-2 on the human-annotated samples, enabling automated labeling. This significantly reduces human effort and streamlines the demonstration adaptor process, making it nearly fully automated. ControlNet Training. We train a ControlNet [controlnet] based on the Stable Diffusion 1.5 [stable_diffusion] model to generate photo-realistic robot images from rendered robot images. To collect training samples, we use teleoperation to gather a small amount of play data, where the robot is teleoperated to move randomly within an empty workspace while recording RGB-D images and corresponding joint states. This ensures a diverse dataset of robot arm configurations, free from occlusions or distractions. Notably, these training samples are platform-specific but task-invariant, meaning they only need to be collected once per robot platform, and the trained ControlNet can be used across all tasks. This also opens up the possibility of directly transforming our in-the-wild demonstrations to other robotic arms without the need to design new exoskeletons that match specific robots. For training, we use a batch size of 88 and a learning rate of , while keeping other hyperparameters at their default settings. We use 50 DDPM sampling steps [ddpm] with a guidance scale of 9.0 [cfg]. The prompt for generating robot images for our robot platform is set to: robotic arms, dual arm, industrial robotic manipulator, metallic silver color, mechanical joints, precise mechanical details, gripper end effector, high-quality photo, photorealistic, clear and sharp details
[…]
[width=]images/dino_feat.pdf
Data Processing. The color image is resized to 448252 for DINOv2 backbone [dino] and 640360 for ResNet-18 backbone [resnet]. The depth image is resized to 640360 before creating the point cloud. The camera intrinsics are adjusted accordingly. Both the point clouds and actions are in the camera coordinate system. The point cloud is down-sampled with a voxel size of 5mm. For the data collected with teleoperation, we crop the point clouds using the range of [-0.7m, 0.7m], [-0.3m, 0.55m] and [0.9m, 1.55m]. For the data collected with AirExo-2, we crop the point clouds using the range of [-0.7m, 0.7m], [-0.3m, 0.45m] and [0.75m, 1.4m]. The robot trajectories are sampled using differences of translation, rotation and gripper width to remove redundant actions. For the action at two adjacent timesteps, if all the differences are less than the thresholds, only the first action is retained. The threshold for translation and gripper width is 5mm and the rotation threshold is . Network. The sparse encoder adopts a ResNet-like architecture built upon MinkowskiEngine [minkowski]. The dense encoder adopts DINOv2-base [dino] as the 2D backbone with the output channel of 128. In the spatial aligner, we use for feature alignment. The aligned features are fused by shared MLPs with the size of (256, 256, 256), and then fed into another sparse network. The two sparse networks are detailed in Tab. H1.. The transformer in action generator contains 4 blocks, in which we set 512 and 2048. The channel number of the readout token is 512. The diffusion head adopts a CNN implementation [dp] with 100 denoising iterations in training and 20 iterations in inference. The output action horizon used in experiments is 20. Training. RISE-2 is trained on 4 Nvidia A100 GPUs. The batch size is 240, the initial learning rate is 3e-4, and the warmup step is 2000. We employ a cosine scheduler to adjust the learning rate during training. 20% of the color images are augmented using a color jitter with (brightness, contrast, saturation, hue) parameters set to (0.4, 0.4, 0.2, 0.1).
[…]
| 2x mean pooling | ||
[…]
We conduct additional experiments on action representations for the Collect Toys task, comparing relative and absolute action representations. The results in Tab. H1. show that while relative action representation sometimes yields better results, absolute action representation provides more stable performance, particularly in terms of generalization. Therefore, we use absolute action representations throughout our experiments, except for the case study in §LABEL:sec:case-study.
[…]
| absolute | 40.0 | 12.5 | 5.0 | |
[…]
| absolute | 65.0 | 45.0 | 42.5 | |
[…]
To evaluate action accuracy, we designed a special evaluation board with three tracks, as illustrated in Fig. H1.. Each track has fixed holes spaced 2 cm apart. We created custom connectors for both the AirExo-2 and UMI [umi] that fit into these fixed holes, allowing us to collect position data. By sequentially placing the connectors into each fixed hole along the track, we can calculate the relative movement distance between two adjacent fixed holes and compare it with the true value (20 mm) to calculate the error.
[width=0.7]images/track.pdf
[…]
Fig. H1. visualizes the sparse semantic features obtained from the dense encoder by projecting the 2D feature map to 3D form using the reference coordinates. The sparse semantic features are aligned to the input point cloud using weighted spatial interpolation detailed in §LABEL:sec:rise2-arch. Although the 2D feature map output by the dense encoder is in low resolution (3218), we still observe clear and distinguishable continuous feature variations on the aligned features, where the targets at the current step can be easily identified from the entire scene. Such characteristic ensures precise feature fusion in the spatial domain. Additionally, we find that the features from DINOv2 change significantly as the task progresses, enabling the model to clearly understand the global state at the current time.

### v2

@@
-
-[…]
-[…]
+references.bib AirExo-2: Scaling up Generalizable Robotic Imitation Learning with Low-Cost Exoskeletons
+Abstract
+Scaling up robotic imitation learning for real-world applications requires efficient and scalable demonstration collection methods.
+While teleoperation is effective, it depends on costly and inflexible robot platforms.
+In-the-wild demonstrations offer a promising alternative, but existing collection devices have key limitations: handheld setups offer limited observational coverage, and whole-body systems often require fine-tuning with robot data due to domain gaps.
+To address these challenges, we present AirExo-2, a low-cost exoskeleton system for large-scale in-the-wild data collection, along with several adaptors that transform collected data into pseudo-robot demonstrations suitable for policy learning.
+We further introduce RISE-2, a generalizable imitation learning policy that fuses 3D spatial and 2D semantic perception for robust manipulations.
+Experiments show that RISE-2 outperforms prior state-of-the-art methods on both in-domain and generalization evaluations.
+Trained solely on adapted in-the-wild data produced by AirExo-2, the RISE-2 policy achieves comparable performance to the policy trained with teleoperated data, highlighting the effectiveness and potential of AirExo-2 for scalable and generalizable imitation learning.
+Keywords: Robotic Manipulation, Scalable Data Collection, Generalizable Imitation Policy
+1 Introduction
+Scaling up generalizable robotic imitation learning in the real world is key to developing robust policies for practical applications [position_realworld, pi0].
+While teleoperation is widely used for collecting demonstrations, it requires physical robots to record both observations and actions for robot-centric demonstrations, making it costly and difficult to scale.
+To address this, recent work explores alternative, low-cost data sources such as human videos [gen2act, phantom, avid, instant_policy, mimicplay, atm, learning_by_watching] and in-the-wild demonstrations [umi, airexo_v1, dobbe_stickv1, dexcap, demoat].
+These human-centric demonstrations eliminate the need for a robot during collection, reducing costs and improving scalability.
+Human video approaches typically infer actions via hand or object pose estimation [ditto, r+x, instant_policy, im2flow2act], while in-the-wild demonstrations use devices that interface human motion with robot-relevant data during task execution.
+Current in-the-wild demonstration devices fall into two main categories: handheld [umi, rum_stickv2, legato, dobbe_stickv1, demoat] and whole-body devices [arcap, airexo_v1, m2r, dexcap].
+Handheld devices are typically designed with end-effectors identical to those of the robot.
+By equipping both the handheld devices and the robot with in-hand cameras, these methods leverage in-hand observations to ensure visual consistency.
+However, they face two key issues: (1) reliance on visual SLAM for pose estimation leads to action inaccuracies (§4.3); and (2) limited camera field of view hinders depth perception during interactions (Appendix C.3).
+On the contrary, whole-body devices offer more accurate action capture and flexible observations, but they typically require fine-tuning with teleoperated data [airexo_v1, dexcap], as in-the-wild demonstrations still exhibit a domain gap compared to real robot data.
+To address these challenges, we present the AirExo-2 system for large-scale in-the-wild demonstration collection and adaptation.
+As illustrated in Fig.
+LABEL:fig:teaser (left), we introduce demonstration adaptors that align both observations and actions from in-the-wild demonstrations with the robot domain, enabling direct use in imitation learning.
+On the hardware side, we build upon AirExo [airexo_v1] with a redesigned exoskeleton featuring a stronger mechanical structure and improved calibration, enhancing action precision and system robustness.
+We conduct comprehensive analyses to evaluate the system’s efficiency and accuracy in demonstration collection.
+Beyond scaling up in-the-wild demonstration collection, developing a robust policy is key to fully leveraging these demonstrations.
+We introduce RISE-2, a generalizable policy that fuses 2D and 3D perception, as illustrated in Fig.
+LABEL:fig:teaser (right).
+RISE-2 outperforms prior imitation policies in both in-domain and generalization settings.
+Notably, it matches the performance of teleoperated-data-trained policies using only in-the-wild demonstrations collected and adapted by AirExo-2 — without any robot data.
+Under equivalent data collection time, it even surpasses teleoperated baselines.
+These results highlight the strong potential of combining AirExo-2 and RISE-2 for scalable, generalizable imitation learning.
+The hardware and codebase will be open-sourced.
+2 Related Works
+Scaling up Demonstration Collection.
+Demonstration data is fundamental to advancing generalizable imitation learning in robotic manipulation.
+Recent studies on data scaling laws [data_scaling_law] reveal that imitation learning in robotics follows similar scaling trends observed in natural language processing [scaling_nlp] and computer vision [scaling_video, dit] fields, highlighting the importance of scaling up demonstration collection.
+Teleoperation is the most direct and widely used approach for collecting real-world robot demonstrations, as it captures data in the robot’s native observation domain and action space.
+However, scaling up teleoperated demonstration collection is challenging, primarily limited by the high cost and infrastructure demands of deploying more robots.
+Consequently, constructing large-scale teleoperated robotic datasets [rt1, agibotworld, oxe, rh20t, bcz, droid, bridgedata, robomind] requires substantial human labor and physical resources.
+Moreover, teleoperation is often less intuitive than natural human manipulation, resulting in higher user learning costs [hajl] and lower demonstration quality [s2i].
+To overcome these limitations, recent research explores generating [gensim2, robotwin, gensim, robogen] or augmenting [rocoda, intervengen, dexmimicgen, mimicgen, cyberdemo] demonstrations automatically in simulation, which facilitates demonstration collection but often necessitates sim-to-real adaptation for real-world deployment.
+Other approaches leverage internet-scale human videos for representation learning [vrb, liv, vip, vc1, r3m, mvp, hrp, mpi] or policy pre-training [gen2act, gr2, vpdd, dexmv, gr1, lapa], yet accurately converting 2D human videos into robot demonstrations with 3D trajectories and robot states remains a challenge [human_video_survey].
+Alternatively, in-the-wild data collection uses specialized human-operated devices — such as hand-held grippers [umi, rum_stickv2, legato, dobbe_stickv1, demoat], cameras [ar2d2, eve], VR/AR systems [arcap, egomimic, m2r], motion-capture gloves [dexcap], and exoskeletons [airexo_v1, m2r] — to translate human motions into robot actions.
+Learning from In-the-Wild Demonstrations.
+Despite promising in scalability, the kinematic gap and the visual gap pose obstacles in learning from in-the-wild demonstrations [airexo_v1, m2r].
+(1) The kinematic gap refers to discrepancies in motion translation between humans and robots, which can lead to inaccuracies in action execution.
+This gap can be addressed through visual or mechanical methods.
+Visual approaches typically use off-the-shelf pose estimation frameworks from commercial devices [arcap, umi, ar2d2, rum_stickv2, egomimic, legato, dobbe_stickv1, dexcap, eve] or structure-from-motion [sfm, demoat] to extract robot poses from demonstrations.
+In contrast, mechanical approaches [airexo_v1, m2r] employ kinematically isomorphic devices that capture robot poses through angle encoder readings.
+(2) The visual gap, on the other hand, arises from the presence of specialized devices and human hands in the captured visual data, whereas robot demonstrations must accurately reflect the robot itself.
+Hand-held devices [umi, rum_stickv2, legato, dobbe_stickv1, demoat] mitigate this by using in-hand cameras and the same end-effector during deployment to maintain visual consistency.
+Other methods address the visual gap by incorporating robot demonstrations for fine-tuning or co-training [ar2d2, airexo_v1, egomimic, eve], or by employing human-in-the-loop strategies to collect corrective behaviors during policy deployment [arcap, dexcap].
+Generalizable Manipulation Policy.
+Direct learning from in-the-wild demonstrations highlights the need for a generalizable manipulation policy capable of transferring learned skills to the robot during real-world deployment.
+A generalizable policy can adapt to new domains or environments, even with limited training data from restricted domains [cage], and should generalize across variations in camera perspectives, backgrounds, objects, and embodiments.
+While many behavior-cloning-based robotic manipulation policies [bc, dp, act3d, rvt, peract, act] perform well in-domain, they often fail in out-of-distribution scenarios, leading to compounded errors and task failures [rise, cage].
+Large-scale pre-training on real-world robot demonstration data can improve generalization [mt-act, rt1, oxe, bcz, openvla, rdt1b, octo, hpt, rt2], but it does not fully address the inherent limit of the policy in adapting across diverse domains.
+Two promising approaches for improving generalization involve leveraging 3D perceptions [polarnet, act3d, rvt, rise, dp3] to enhance environmental understanding and using 2D foundation models [what_pretrain, spawnnet, soft, cage, same] to enrich the policy’s recognition of complex object and scene information.
+Recent work [lift3d, sgr, sgrv2] combines these approaches to further enhance policy performance, though challenges remain in effectively integrating these strategies for broader generalization across diverse manipulation domains.
+3 Method
+3.1 AirExo-2: Efficiently Collecting and Adapting In-the-Wild Demonstrations
+As discussed in §2, eliminating the need for physical robots enables low-cost and scalable in-the-wild demonstration collection.
+However, visual and kinematic discrepancies introduce significant challenges for directly utilizing such data.
+To overcome this, our objective is to efficiently convert in-the-wild demonstrations into pseudo-robot demonstrations suitable for training real-world manipulation policies.
+This requires (1) robust hardware for reliable data collection, (2) accurate motion tracking, and (3) photo-realistic visual adaptation to support effective policy learning.
+To this end, we present the AirExo-2 system, designed to efficiently collect and adapt in-the-wild demonstrations for downstream manipulation policy training, as illustrated in Fig.
+1.
+Unlike prior approaches that circumvent the visual gap through in-hand cameras [umi, rum_stickv2, legato, dobbe_stickv1, demoat] or rely on pre-training and fine-tuning paradigm [arcap, ar2d2, airexo_v1, egomimic, dexcap, eve], our system explicitly bridges this gap by converting in-the-wild observations into pseudo-robot views using visual adaptors.
+Combined with accurate motion tracking enabled by the mechanical design, this semi-automatic adaptation process better aligns demonstration data with robot deployment conditions, allowing policies to directly transfer to real-world robot platform.
+Please refer to Appendix A for implementation details.
+Hardware.
+As shown in Fig.
+1 (left), we upgrade the AirExo design [airexo_v1] with aluminum profiles to improve structural robustness and motion tracking precision.
+To facilitate subsequent demonstration adaptations, we design the AirExo-2 to have a 1:1 scale and maintain kinematic isomorphism with the robotic arm.
+A mobile base is added to support data collection in diverse environments, and a dedicated calibration process is introduced to ensure accurate motion capture.
+Details of the hardware design and calibration procedure are provided in Appendix A.1 and Appendix A.2, respectively.
+The system retains the low-cost advantage of the original AirExo, with the dual-arm platform (excluding the camera) priced at just $600.
+Operation Space Adaptor.
+Following [rise, idp3], we project all states and actions into the global camera coordinate — a universal coordinate frame between the AirExo-2 and robot platform.
+Image Adaptor.
+As shown in Fig.
+1 (top), we convert in-the-wild image into pseudo-robot image by (1) inpainting human-related regions — including hands and AirExo-2 — and (2) overlaying it with the generated robot image.
+Leveraging the joint encoder readings, calibration results, and the joint mapping between AirExo-2 and the robot, we render RGB-D and mask images of both systems in Open3D [open3d].
+The rendered AirExo-2 mask is merged with the hand mask obtained via SAM-2 [sam2] to localize human-related areas, which are then inpainted using ProPainter [propainter].
+Inspired by [rovi_aug], the rendered robot image is refined via ControlNet-guided Stable Diffusion [stable_diffusion, controlnet] and composited onto the inpainted image to generate the final photo-realistic pseudo-robot image.
+Depth Adaptor.
+As shown in Fig.
+1 (bottom), we first capture a universal reference depth of the empty workspace, and combine it with the static object depth from the first frame to form a reference depth for each demonstration.
+Guided by the combined mask from the image adaptor, we inpaint human-related regions in the depth map with the reference depth to maintain spatial consistency.
+The adapted depth is obtained by merging the inpainted depth with the rendered robot depth.
+3.2 RISE-2: A Generalizable Policy for Learning from In-the-Wild Demonstrations
+Although in-the-wild demonstrations are adapted into pseudo-robot demonstrations, domain gaps — such as differences in camera viewpoints — still persist.
+To effectively learn from such data, a generalizable policy is required.
+As discussed in §2, 3D perception provides view-invariant geometric information, while 2D foundation models offer rich semantic understanding from large-scale pretraining.
+To leverage the complementary strengths of both, we introduce RISE-2, a generalizable policy that integrates 3D spatial perception with 2D semantic features for robust robotic manipulation.
+As illustrated in Fig.
+2, the RISE-2 policy consists of a sparse encoder for 3D features, a dense encoder for 2D features, a spatial aligner for feature fusion, and an action generator that maps fused features to actions.
+Implementation details are provided in Appendix B.
+Sparse Encoder.
+3D point cloud data contains rich spatial structure, enabling effective extraction of local geometric features — a property widely used in general grasping [anygrasp, fang2020graspnet, wang2021graspness].
+RISE [rise] encodes point clouds using raw color information to capture both semantic and geometric cues, but it struggles to disentangle coordinate shifts from color variations.
+RISE-2 inherits the sparse 3D encoder from RISE, but removes color input to focus purely on geometric features.
+Formally, the sparse encoder , maps a depth image and camera intrinsics to sparse geometric features , along with the corresponding seed point coordinates after network down-sampling.
+The encoder adopts a ResNet-like architecture [resnet] built on MinkowskiEngine [minkowski], enabling efficient sparse feature extraction and supporting real-time policy execution.
+Dense Encoder.
+A generalizable policy requires rich semantic features to understand the scene, while the low-quality texture of point cloud data poses a challenge to this demand.
+Hence, RISE-2 adopts a dense 2D encoder to focus on semantic understanding based on the color image.
+Formally, the dense encoder processes the color image into semantic feature map densely organized in the 2D form of shape , along with its reference 3D coordinates , calculated by applying 2D adaptive average pooling to size on the point cloud generated from the depth map and camera intrinsics .
+One key advantage of the dense encoder is its compatibility with 2D foundation models, which provides highly generalizable visual representations across diverse tasks and domains [cage].
+We implement using DINOv2 [dino] fine-tuned with LoRA [lora], enhancing the robustness and adaptability of the policy in capturing contextual relationships within the environment.
+Spatial Aligner.
+RISE-2 extracts geometric and semantic features using separate encoders, posing a challenge for feature fusion from different domains.
+A straightforward solution is to concatenate the aggregated feature vectors, but this discards fine-grained local information critical for precise perception.
+An alternative [sgr, sgrv2] upsamples to match the resolution of the input image , projects it onto the point cloud, and then downsamples to align with the seed points .
+However, this process is computationally expensive and reduces the efficiency of policy training and deployment.
+Instead, RISE-2 utilizes a spatial aligner to efficiently fuse two kinds of features based on their 3D coordinates and .
+For each point output by the sparse encoder , we compute its nearest neighbors from the coordinates output by the dense encoder .
+The aligned semantic feature of is computed by weighted spatial interpolation:
+where denotes the Euclidean distance between and .
+Then, the aligned feature for point is obtained by concatenation, i.e., .
+By aligning the seed points with the dense semantic feature map through their 3D coordinates, we can accurately retrieve fine-grained semantic features at arbitrary locations.
+The fused features are then aggregated into high-level sparse representations using sparse convolution layers [minkowski].
+Visualization of the weighted spatial interpolation applied to the 2D feature map is provided in Appendix B.3.
+Action Generator.
+RISE-2 utilizes a decoder-only transformer [vaswani2017attention] to map point tokens with sparse positional encoding [rise] to the latent feature, which conditions a diffusion head [dp, ddpm, janner2022planning] to generate the action chunk [act].
+Actions are represented in the camera frame for cross-scene consistency, with translations in absolute coordinates and rotations in the 6D format [rot6d].
+4 Experiments
+4.1 Policy Evaluation: RISE-2
+Setup.
+To evaluate the performance of policies in both in-domain and generalization settings, we design 4 tasks in Fig.
+3 and collect 50 demonstrations for each task using teleoperation [airexo_v1].
+We compare RISE-2 against several representative policies based on 2D images or 3D point clouds, including ACT [act], Diffusion Policy [dp], CAGE [cage], RISE [rise], and [pi0].
+Additionally, we include a RISE-2 variant with the dense encoder implemented with an ImageNet-pretrained [imagenet] ResNet-18 [resnet], denoted as RISE-2 (ResNet-18).
+Please refer to Appendix C.1 for details.
+In-Domain Evaluation.
+As shown in Fig.
+3, RISE-2 significantly outperforms previous state-of-the-art imitation policies in average success rate across 4 tasks, demonstrating its effectiveness in handling diverse manipulation challenges.
+On tasks requiring precise motion execution, such as Lift Plate and Open Lid, RISE-2 achieves superior accuracy in predicting fine-grained actions, consistently surpassing all baselines.
+The performance gain over RISE [rise] highlights the benefits of using separate 2D and 3D encoders with spatial feature alignment.
+Further improvements from RISE-2 (ResNet-18) to RISE-2 (DINOv2) underscore the value of leveraging diverse knowledge from visual foundation models [dino] in robotic manipulation.
+| Method | In-Domain | Generalization | ||
+| Bg. | Obj. | Both | ||
+| ACT [act] | 32.5% | 0.0% | 2.5% | 0.0% |
+| Diffusion Policy [dp] | 40.0% | 12.5% | 5.0% | 2.5% |
+| CAGE [cage] | 65.0% | 45.0% | 42.5% | 10.0% |
+| RISE [rise] | 72.5% | 42.5% | 12.5% | 40.0% |
+| [pi0] | 85.0% | 77.5% | 82.5% | 82.5% |
+| RISE-2 (ResNet-18) | 77.5% | 47.5% | 47.5% | 37.5% |
+| RISE-2 (DINOv2) | 95.0% | 85.0% | 85.0% | 62.5% |
+Generalization Evaluation.
+We select the Collect Toys task to evaluate the generalization ability of different policies under varying environmental disturbances.
+Please refer to Appendix C.1 for disturbance details.
+As shown in Tab.
+1, RISE-2 demonstrates strong robustness, maintaining high performance even under combined disturbances.
+While the RISE-2 (ResNet-18) variant performs slightly below the RISE-2 (DINOv2) model, it still generalizes well and outperforms most baselines.
+These results further validate our architectural choice of using separate encoders for 3D geometric and 2D semantic features, which effectively enhances policy robustness.
+4.2 System Evaluation: AirExo-2
+Learning from Pseudo-Robot Demonstrations.
+Previous experiments have demonstrated that RISE-2 is a generalizable policy, making it well-suited for learning from in-the-wild demonstrations collected and processed by AirExo-2.
+In this evaluation, we utilize AirExo-2 to collect an equal number of in-the-wild demonstrations and convert them into pseudo-robot demonstrations.
+These demonstrations are then used to train the RISE-2 policy, which is deployed on a real robot without further fine-tuning.
+As shown in Fig.
+4, the policy achieves satisfactory success rates on all tasks, with only a slight performance drop compared to the policy trained on teleoperated demonstrations.
+This highlights the ability of the AirExo-2 system to collect and adapt high-quality demonstrations suitable for training downstream imitation policies.
+Please refer to Appendix C.4 for additional results of the RISE policy [rise].
+Complex Task.
+We evaluate the AirExo-2 system and the RISE-2 policy on the challenging, long-horizon, and contact-rich task Serve Steak.
+Using 50 in-the-wild demonstrations collected and adapted by AirExo-2, the RISE-2 policy achieves a decent 50% overall success rate, with strong performance even in the contact-rich Scoop Steak phase, as shown in Fig.
+5.
+This demonstrates the effectiveness of AirExo-2 in capturing accurate, fine-grained actions and the ability of RISE-2 to accomplish complex tasks without requiring robot data.
+Combined with the scalability analysis in §4.3, increasing the number of in-the-wild demonstrations is expected to further boost performance.
+These results highlight the strong synergy between scalable, high-quality demonstration collection by AirExo-2 and generalizable RISE-2 policy learning, paving the way for efficient and adaptable robotic systems in real-world environments.
+| Method | Success Rate | |
+| w.o. adaptors | w. adaptors | |
+| RISE [rise] | 30.0% | 57.5% |
+| RISE-2 | 52.5% | 90.0% |
+Ablation on Visual Adaptors.
+We use the Collect Toys task to highlight the importance of the demonstration adaptor for learning from in-the-wild data.
+As shown in Tab.
+2, removing visual adaptors significantly reduces performance, revealing a substantial visual gap between in-the-wild and robot demonstrations.
+These results also validate the quality of AirExo-2 pseudo-robot demonstrations for training high-performing policies.
+Notably, RISE-2 without adaptors still performs comparably to RISE [rise] with adaptors, demonstrating its strong generalization across embodiments.
+4.3 System Analysis: AirExo-2
+| Method | Completion Time (s) | Avg. Rank | Rating | |
+| Intui. | Learn. | |||
+| EE Teleop. | 3.00/3 | 2.95/3 | 29.75 | |
+| Joint Teleop. | 1.80/3 | 2.00/3 | 49.58 | |
+| AirExo-2 | 1.20/3 | 1.05/3 | 83.00 | |
+User Study.
+We compare three demonstration collection methods: end-effector pose teleoperation using a haptic device [rh20t], joint-space teleoperation via AirExo [airexo_v1], and in-the-wild collection with AirExo-2.
+Twenty participants (14 men, 6 women, aged 21–35) of varying expertise levels perform the Collect Toys task on all three platforms in randomized order, with each trial consisting of a 3-minute familiarization and a timed demonstration collection.
+After completing all trials, participants fill out a questionnaire to provide feedback on three collection methods.
+From the results in Tab.
+3, we observe that both experienced and novice users find AirExo-2 easier to learn and operate than both types of teleoperation, leading to faster onboarding and smoother operation.
+Statistical analysis (Welch’s -test, ) further confirms its superior efficiency in task completion time.
+The user-friendliness and efficiency make AirExo-2 well-suited for diverse users and consistent, high-quality data collection at scale in real-world settings.
+| Device | Error (mm) | |
+| Avg | Max | |
+| UMI [umi] | 20.002 | |
+| AirExo-2 | 6.134 | |
+Action Accuracy.
+Accurate actions are essential for learning reliable policies from in-the-wild demonstrations.
+Unlike handheld devices such as UMI [umi], which use visual SLAM for trajectory estimation, AirExo-2 leverages its mechanical structure and forward kinematics for more precise motion tracking.
+As shown in Tab.
+4, it achieves significantly lower translation errors across 3 tracks (please refer to Appendix C.2 for details about accuracy evaluations), confirming its reliability for high-fidelity motion capture.
+Scalability Analysis.
+We compare the scalability and effectiveness of AirExo-2 in-the-wild data collection with traditional teleoperation data collection in Fig.
+6.
+Given the same collection time, AirExo-2 yields higher demonstration throughput (analogous to the findings in the previous user study) with lower costs ($0.6k per AirExo-2 system versus approximately $60k per teleoperation platform with robots), making it a more scalable solution to demonstration collection.
+Moreover, using the demonstrations collected under the same time, policies trained on AirExo-2 demonstrations outperform those trained on teleoperation data, underscoring the value of adapted pseudo-robot demonstrations for efficient and effective imitation learning.
+5 Conclusion
+This paper introduces AirExo-2, a novel system designed for scalable in-the-wild demonstration collection and adaptation using low-cost exoskeletons.
+By incorporating several adaptors, AirExo-2 enables the visual and kinematic transformation of in-the-wild demonstrations into pseudo-robot demonstrations, which can then be directly applied to downstream imitation learning tasks.
+We also propose a generalizable policy, RISE-2, which effectively integrates both 2D and 3D perception, demonstrating exceptional performance in both in-domain and generalization scenarios.
+Further experiments on a variety of tasks demonstrate that when trained exclusively on pseudo-robot demonstrations generated by the AirExo-2 system — without using any robot data — the policy achieves satisfactory performance during zero-shot deployment on a real-world robot platform, even for the complex, long-horizon, and contact-rich task Serve Steak.
+This highlights the potential of combining AirExo-2 and RISE-2 as a scalable and promising alternative to traditional teleoperation-imitation pipelines, providing a more efficient, cost-effective solution for scalable, generalizable robotic imitation learning.
+Together, these results open new possibilities for transferring manipulation skills from in-the-wild environments to real robots, without the need for extensive robot-centric data collection.
+6 Limitations and Future Works
+While we utilize the proposed demonstration adaptor to visually transform in-the-wild demonstrations collected by AirExo-2 into pseudo-robot demonstrations, these transformed demonstrations are primarily useful for generalizable policies.
+To enhance the applicability of the pseudo-robot demonstrations to a broader range of policies, future work could explore the integration of demonstration augmentation methods, such as novel view synthesis [rovi_aug, zeronvs, vista, gcd], into the adaptation process to improve the diversity of the demonstrations, making them more versatile for various policy learning.
+As demonstrated by several works [umi, inhand_than_global, giving_robots_a_hand, demoat], the in-hand image is a semi-unified observation modality across different embodiments.
+Our case study also reveals that combining in-hand observations can strengthen the performance of various 2D policies.
+However, the current AirExo-2 system does not include in-hand cameras.
+Although we have designed connectors to integrate them with the exoskeleton, calibrating the in-hand cameras with the exoskeleton remains challenging, making it difficult to adapt the in-hand images into the robot domain.
+Future work could explore effective methods for adapting in-hand images collected by AirExo-2 to the robot domain, or investigate strategies for leveraging the semi-unified in-hand observations in policy design.
+The current AirExo-2 system only supports parallel grippers as end-effectors, limiting its applicability in more dexterous tasks.
+Future work could integrate the AirExo-2 system with dexterous hands [eyesight_hand, leap_hand] and their corresponding exoskeletons [ben2025homie, other_hand], enabling more complex manipulation capabilities and expanding the range of tasks the system can effectively perform.
+Moreover, adding additional sensors to the AirExo-2 system for recording other observation modalities — such as force/torque [foar, forcemimic], tactile [vitamin], and sound [maniwav] — can also be explored in the future.
+Appendix
+Appendix A AirExo-2 Implementation
+A.1 Hardware Design
+High-quality in-the-wild data requires precise collection for effective pseudo-robot transformation. However, AirExo [airexo_v1] has several key hardware limitations:
 H1.
@@
 H5.
-Its shaft connects directly to the encoder, with wires routed along its side. Joint movement can cause constant friction and stretching, risking wire breakage or short circuits over long-time use.
+The shaft connects directly to the encoder, with side-routed wires that endure friction and stretching during joint movement, risking long-term damage or short circuits.
 During demonstration collection using teleoperation, most of the above issues (except H4) might not significantly impact data acquisition, as the human operator can adjust their actions based on the movements of the robotic arms.
 However, during the in-the-wild demonstration collection, these drawbacks can substantially affect motion capture accuracy and may result in invalid or unusable demonstration data.
-The hardware design of AirExo-2 is primarily driven by the limitations mentioned above.
-To ensure seamless integration with the learning process from in-the-wild demonstrations, we design the exoskeleton to match the dimensions of the robotic arm in a 1:1 ratio.
-This design choice helps avoid unnecessary obstacles in direct learning from the collected demonstrations.
-The key hardware designs are outlined as follows.
+Driven by the limitations above, the hardware design of AirExo-2 focuses on seamless integration with the learning process from in-the-wild demonstrations.
+To enable direct learning, the exoskeleton is built with a 1:1 dimensional match to the robotic arm, minimizing discrepancies between human and robot embodiments.
+An overview of the hardware is shown in Fig.
+1.
+Excluding cameras, the complete AirExo-2 platform costs $600.
+Key hardware design elements are outlined below.
 Enhanced Overall Structural Rigidity (H1).
@@
 Inside the joint, larger bearings are used to further enhance structural rigidity.
-Together with the improved links, the hardware upgrade significantly increases the overall structural rigidity of AirExo-2, making the exoskeleton more durable, and thereby improving data collection accuracy.
+Together with the improved links, the hardware upgrade significantly increases the overall structural rigidity of AirExo-2, making the exoskeleton more durable, and thereby improving system accuracy.
 Hollow Rotating Disc and Side-Mounted Encoder (H5).
 As shown in Fig.
-H1., this design features an encoder mounted on the side of the joint, which uses a gear mechanism to translate the rotational angle of the rotating disc into encoder readings.
+7, this design features an encoder mounted on the side of the joint, which uses a gear mechanism to translate the rotational angle of the rotating disc into encoder readings.
 The hollow disc design allows wires to pass through, preventing them from stretching during rotation and thereby extending their lifespan.
@@
 As illustrated in Fig.
-H1., the rotating disc, pre-joint, and post-joint are connected through bearings, allowing the rotational motion of the joint to be directly transmitted to the track.
+7, the rotating disc, pre-joint, and post-joint are connected through bearings, allowing the rotational motion of the joint to be directly transmitted to the track.
 For demonstration collection, excessive or insufficient friction in the rotation of the joints is undesirable.
@@
 This design ensures optimal friction for comfortable and accurate data collection.
-[width=]images/hardware.pdf
 Smooth Gripper Control (H4).
 Following [umi, forcemimic], the gripper of AirExo-2 incorporates a linear guide, with the fingers mounted on a sliding block.
-This design allows for smoother opening and closing of the gripper, ensuring it operates seamlessly even under significant clamping forces without any stalling.
+This design allows for smoother opening and closing of the gripper, ensuring it operates seamlessly under clamping forces without any stalling.
 Mobile Data Collection Platform (H3).
@@
 However, to address the issue of base movement caused by the body motion of the operator, we mount AirExo-2 on a mobile aluminum profile stand, as shown in Fig.
-H1..
+7.
 This setup ensures stability of the base during demonstration collection while maintaining the flexibility needed for mobility, enabling large-scale demonstration collection in real-world environments.
@@
 We also designed two optional camera mounts (though not used in this paper) for the future integration of in-hand cameras on the top of both grippers.
-[…]
+A.2 Calibration
 The AirExo-2 system requires two types of calibration simultaneously: (1) aligning the zero positions of each joint with the corresponding robot joint, and (2) determining the transformation between the global camera and the AirExo-2 base.
@@
 For initial calibration, the former calibration can be achieved by manually adjusting the joints to approximate the zero position using specialized 3D-printed tools and reading the encoder values, obtaining .
-The latter calibration can be done by attaching an ArUco calibration marker board with a known position on the base and performing optical calibration using the OpenCV library [opencv], obtaining .
+The latter calibration can be done by attaching an ArUco calibration marker board with a known position on the base and performing optical calibration using OpenCV [opencv], obtaining .
 Thus, the transformation can calculated as
@@
 This iterative refinement compensates for errors that accumulate across joints, ultimately improving the overall accuracy.
-Specifically, we define , the calibration parameters to be optimized through differentiable rendering, as
+Specifically, we define , the calibration parameters to be optimized, as
 where parameters, except for the base-to-camera rotation, are represented as deltas relative to the initial calibration results.
@@
 Thus, the final calibration results can be calculated as
-{gather}
-[…]
-[…]
 and the initial parameter values are set as
@@
 Subsequently, we utilize SAM-2 [sam2] to annotate the AirExo-2 mask and the depth mask for each sample , as shown in Fig.
-H1..
+8.
 The first mask provides supervision for the rendered AirExo-2 mask, while the second mask is used to select the valid depth information that serves as the supervision signal.
-[width=]images/calib_detail.pdf
 The differentiable rendering engine Redner [redner] is employed to render the AirExo-2 mask and AirExo-2 depth using the calibration results and joint information:
@@
 In practice, we set , use samples for optimization, and employ the Adam optimizer [adam] with a learning rate of for 1000 iterations to fine-tune the calibration results.
-[…]
-Apart from the two-stage calibration process described in Appendix H1., we implement several calibration alternatives, including (1) initial calibration, which uses the calibration results from the first stage without further fine-tuning; (2) human annotation, where a human operator utilizes a real-time GUI program to manually adjust the calibration parameters to align the rendered AirExo-2 contour with the visually observed contour from camera frames; and (3) two-stage calibration (mask only), which fine-tunes the calibration results by using only mask differences as supervision.
-[…]
-[…]
-[…]
+| Calibration Method | Difference | |
+| Mask (%) | Depth (mm) | |
+| Initial Calibration | ||
 | Human Annotation | ||
@@
 | Two-Stage Calibration (mask + depth) | ||
-[…]
-Our two-stage calibration process achieves more accurate results compared to other alternatives.
-As reported in Tab.
-H1., our two-stage calibration process achieves the lowest error rates, with a 0.78% mask difference and 14.0 mm depth difference, yielding more precise calibration.
-It’s worth noticing that the depth difference here does not fully represent action accuracy because commercial depth sensors can produce noisy depth maps.
-Please refer to §LABEL:sec:accuracy-anaylsis for more details about action accuracy.
-Interestingly, human annotation performs even worse than initial calibration, mainly because annotators can only rely on 2D visual information to adjust calibration parameters.
-This limitation makes it difficult to accurately estimate depth information, leading to larger errors.
-Conversely, our two-stage calibration process explicitly models 3D information via differentiable rendering, offering a more reliable and precise solution for calibrating the AirExo-2 system.
-[…]
+Calibration Analysis.
+Precise calibration is vital for accurate action tracking.
+In addition to our proposed full two-stage calibration method, we evaluate three alternatives: (1) initial calibration, using first-stage results without refinement; (2) human annotation, where a user manually adjusts calibration parameters via a real-time GUI; and (3) two-stage calibration (mask only), which refines calibration using only mask differences as supervision.
+As shown in Tab.
+5, our proposed full two-stage method achieves the highest accuracy, with a 0.78% mask difference and 14.0 mm depth error.
+While depth errors are affected by sensor noise and cannot fully reflect action accuracy (please refer to §4.3 for action accuracy analysis), the results confirm the benefits of modeling 3D information through differentiable rendering for accurate calibration.
+Notably, human annotation even performs worse than initial calibration due to its reliance on 2D visual cues, which limits depth estimation accuracy.
+A.3 Visual Adaptors
 Semi-Automatic SAM-2 Annotations.
 In the image adaptor, we initially annotate the hand mask (and, if visible, the head mask) manually using SAM-2 [sam2].
-However, after a few annotations, we can fine-tune SAM-2 on the human-annotated samples, enabling automated labeling.
-This significantly reduces human effort and streamlines the demonstration adaptor process, making it nearly fully automated.
+After a few annotations on approximately 50 scenes, we can fine-tune SAM-2 on human-annotated samples, enabling automated labeling.
+This significantly reduces human effort and streamlines the demonstration adaptation process, making it nearly fully automated.
 ControlNet Training.
@@
 We use 50 DDPM sampling steps [ddpm] with a guidance scale of 9.0 [cfg].
-The prompt for generating robot images for our robot platform is set to: robotic arms, dual arm, industrial robotic manipulator, metallic silver color, mechanical joints, precise mechanical details, gripper end effector, high-quality photo, photorealistic, clear and sharp details
-[…]
-[width=]images/dino_feat.pdf
-Data Processing.
+The prompt for generating robot images for our robot platform is displayed as follows.
+Appendix B RISE-2 Implementation
+B.1 Data Processing
+Image.
 The color image is resized to 448252 for DINOv2 backbone [dino] and 640360 for ResNet-18 backbone [resnet].
@@
 The camera intrinsics are adjusted accordingly.
-Both the point clouds and actions are in the camera coordinate system.
-The point cloud is down-sampled with a voxel size of 5mm.
-For the data collected with teleoperation, we crop the point clouds using the range of [-0.7m, 0.7m], [-0.3m, 0.55m] and [0.9m, 1.55m].
-For the data collected with AirExo-2, we crop the point clouds using the range of [-0.7m, 0.7m], [-0.3m, 0.45m] and [0.75m, 1.4m].
-The robot trajectories are sampled using differences of translation, rotation and gripper width to remove redundant actions.
+Both the point clouds and actions are projected to the camera coordinate system following [rise, idp3].
+Point Cloud.
+The point cloud is down-sampled with a voxel size of 5mm. For the data collected with teleoperation, we crop the point clouds using the range of
+For the data collected with AirExo-2, we crop the point clouds using the range of
+Robot Trajectory Sampling.
+The robot trajectories are sampled using differences of translation, rotation, and gripper width to remove redundant actions.
 For the action at two adjacent timesteps, if all the differences are less than the thresholds, only the first action is retained.
-The threshold for translation and gripper width is 5mm and the rotation threshold is .
-Network.
+The threshold for translation and gripper width is 5mm, and the rotation threshold is .
+B.2 Hyperparameters
+Policy.
 The sparse encoder adopts a ResNet-like architecture built upon MinkowskiEngine [minkowski].
@@
 The two sparse networks are detailed in Tab.
-H1..
-The transformer in action generator contains 4 blocks, in which we set 512 and 2048.
+6.
+The transformer in action generator contains 4 blocks, in which we set and following [rise].
 The channel number of the readout token is 512.
@@
 The output action horizon used in experiments is 20.
+| Layer Name | Sparse Encoder | Spatial Aligner |
+| init_conv | , , | - |
+| 2x mean pooling | ||
+| conv1 | , , | , , |
+| conv2 | , , | , , |
+| conv3 | , , | , , |
+| conv4 | , , | , , |
+| final_conv | , , | , , |
 Training.
-RISE-2 is trained on 4 Nvidia A100 GPUs.
-The batch size is 240, the initial learning rate is 3e-4, and the warmup step is 2000.
+We train the RISE-2 policy on 4 NVIDIA A100 GPUs for 60k steps with a batch size of 240, an initial learning rate of , and a warmup step of .
 We employ a cosine scheduler to adjust the learning rate during training.
-20% of the color images are augmented using a color jitter with (brightness, contrast, saturation, hue) parameters set to (0.4, 0.4, 0.2, 0.1).
-[…]
-[…]
-| 2x mean pooling | ||
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-We conduct additional experiments on action representations for the Collect Toys task, comparing relative and absolute action representations.
-The results in Tab.
-H1.
-show that while relative action representation sometimes yields better results, absolute action representation provides more stable performance, particularly in terms of generalization.
-Therefore, we use absolute action representations throughout our experiments, except for the case study in §LABEL:sec:case-study.
-[…]
-[…]
-[…]
-| absolute | 40.0 | 12.5 | 5.0 | |
-[…]
-| absolute | 65.0 | 45.0 | 42.5 | |
-[…]
-[…]
-To evaluate action accuracy, we designed a special evaluation board with three tracks, as illustrated in Fig.
-H1..
+During training, 20% of the color images are augmented using color jitter with (brightness, contrast, saturation, hue) parameters set to (0.4, 0.4, 0.2, 0.1).
+B.3 Visualization of Sparse Semantic Features
+Fig.
+9 visualizes the sparse semantic features obtained from the dense encoder by projecting the 2D feature map to 3D form using the reference coordinates.
+The sparse semantic features are aligned to the input point cloud using weighted spatial interpolation, as detailed in §3.2.
+Although the 2D feature map output by the dense encoder is in low resolution (3218 for DINOv2 [dino] backbone), we still observe clear and distinguishable continuous feature variations on the aligned features, where the targets at the current step can be easily identified from the entire scene.
+Such a characteristic ensures precise feature fusion in the spatial domain.
+Additionally, we find that the features from the visual foundation model DINOv2 [dino] change significantly as the task progresses, enabling the model to clearly understand the global state at the current time.
+Appendix C Experiments Details
+C.1 Setup
+Data Collection.
+The teleoperated demonstrations are collected using AirExo [airexo_v1], while in-the-wild demonstrations are collected and adapted using our proposed AirExo-2 system.
+For each task, unless further specified, we collect 50 teleoperated demonstrations for RISE-2 policy evaluation and 50 in-the-wild demonstrations for AirExo-2 system evaluations.
+Baselines.
+We compare RISE-2 against several representative policies based on 2D images and 3D point clouds, including: (1) ACT [act], which employs transformers to map image observations and proprioception to robot action chunks; (2) Diffusion Policy [dp], which formulates action prediction as a diffusion denoising process [ddpm, ddim] conditioned on the image observations; (3) CAGE [cage], an extension of Diffusion Policy that incorporates visual foundation models [dino], a causal observation perceiver [perceiver], and an attention-based diffusion action head for improved generalization; (4) RISE [rise], a 3D imitation policy that leverages a sparse 3D encoder for efficient point cloud perception; and (5) [pi0], a vision-language-action flow model for general robot control.
+Evaluation Protocols.
+All policies except are deployed on a workstation with an NVIDIA RTX 2060 SUPER GPU, while is deployed remotely on a server with an NVIDIA A100 GPU for GPU memory considerations.
+Following the procedure outlined in [umi, cage], we adopt a consistent evaluation method for each policy to minimize performance variation and ensure reproducibility.
+Specifically, we generate uniformly distributed test positions randomly before each task evaluation.
+The workspace is set up identically across different policies and test environments, and success rates are recorded for each test case.
+Each policy is evaluated over 20 consecutive trials per task (except for the Serve Steak task, where we evaluate each policy for 10 consecutive trials), and the success rates are computed accordingly.
+Generalization Evaluation.
+We select the Collect Toys task to conduct a generalization experiment to evaluate the robustness of different policies under varying levels of environmental disturbances.
+The policies are trained with demonstrations from a narrow domain, with only one training background and one target object.
+Then, we introduce two types of disturbances during generalization evaluations: background variations and object differences, as illustrated in Fig.
+10.
+We then follow the same evaluation protocol introduced above to compute the policy success rates under (1) novel backgrounds, (2) novel objects, and (3) both novel backgrounds and novel objects, enabling a comprehensive comparison of each policy’s generalization capability.
+For evaluation concerning the background variations, we use each of the 5 novel backgrounds for 4 evaluation trials in the same order to ensure fairness.
+C.2 Action Accuracy Analysis Details
+To evaluate the action accuracy of the data collection systems, we designed a special evaluation board with three tracks, as illustrated in Fig.
+11.
 Each track has fixed holes spaced 2 cm apart.
@@
 By sequentially placing the connectors into each fixed hole along the track, we can calculate the relative movement distance between two adjacent fixed holes and compare it with the true value (20 mm) to calculate the error.
-[width=0.7]images/track.pdf
-[…]
-Fig.
-H1.
-visualizes the sparse semantic features obtained from the dense encoder by projecting the 2D feature map to 3D form using the reference coordinates.
-The sparse semantic features are aligned to the input point cloud using weighted spatial interpolation detailed in §LABEL:sec:rise2-arch.
-Although the 2D feature map output by the dense encoder is in low resolution (3218), we still observe clear and distinguishable continuous feature variations on the aligned features, where the targets at the current step can be easily identified from the entire scene.
-Such characteristic ensures precise feature fusion in the spatial domain.
-Additionally, we find that the features from DINOv2 change significantly as the task progresses, enabling the model to clearly understand the global state at the current time.
+C.3 Case Study: Are In-Hand Cameras Sufficient for Manipulations?
+We conducted a case study to investigate whether in-hand cameras are sufficient for many manipulation tasks.
+We select the Collect Toys task as an example and utilize CAGE [cage] as the policy for this case study.
+In-hand cameras alone are often insufficient for manipulation tasks and may pose additional obstacles on policy learning.
+As shown in Fig.
+12 (bottom), neither global nor in-hand cameras alone provide adequate observations for 2D image-based policies to achieve strong performance.
+Recent work [rise] has demonstrated that using only a global camera enables a 3D imitation policy to outperform 2D multi-view image-based policies, highlighting the importance of 3D information for scene understanding.
+However, as illustrated in Fig.
+12 (top), in-hand cameras may produce incomplete depth information when the robotic arm approaches an object, making them unsuitable for 3D point-cloud-based policies.
+Consequently, relying solely on in-hand cameras can degrade the performance of the policies, particularly for 3D policies that rely on complete and accurate depth information to achieve superior performance.
+C.4 Additional Experiment Results
+Learning from Pseudo-Robot Demonstrations.
+We also evaluate the performance of the RISE policy [rise] using pseudo-robot demonstrations.
+Specifically, we employ AirExo-2 to collect an equivalent number of in-the-wild demonstrations, which are then transformed into pseudo-robot data and used to train the RISE policy.
+The trained policy is deployed on a real robot without any fine-tuning.
+As shown in Fig.
+13 and Tab.
+7, RISE achieves good zero-shot performance, though with a slightly larger drop compared to RISE-2.
+These findings highlight the critical role of a generalizable policy in successfully transferring manipulation skills from in-the-wild demonstrations to real robots, under the circumstances of no teleoperation data.
+| Data Source | Policy | Average Success Rate (4 tasks) |
+| Teleoperation | RISE [rise] | 74.375% |
+| RISE-2 | 93.750% | |
+| AirExo-2 | RISE [rise] | 60.625% -13.750% |
+| RISE-2 | 85.000% |
+Action Representations.
+We conduct additional experiments on action representations for the Collect Toys task, comparing relative and absolute action representations.
+The results in Tab.
+8 show that while relative action representation sometimes yields better results, absolute action representation provides more stable performance, particularly in terms of generalization.
+Therefore, we use absolute action representations for Diffusion Policy [dp] and CAGE [cage] throughout our experiments, except for the case study in Appendix C.3.
+For [pi0], we still use relative action representation for consistency with its released pre-trained models.
+| Method | Action Representation | Success Rate | ||
+| in-domain | background | object | ||
+| Diffusion Policy [dp] | relative | 37.5% | 20.0% | 5.0% |
+| absolute | 40.0% | 12.5% | 5.0% | |
+| CAGE [cage] | relative | 72.5% | 32.5% | 35.0% |
+| absolute | 65.0% | 45.0% | 42.5% | |